OpenStudy (anonymous):

needing help with proble, log2(x+1) +log2(x-1)=3

OpenStudy (saifoo.khan):

\[\Large \log_2 [(x+1)(x-1)] = 3\]

OpenStudy (saifoo.khan):

\[\Large [(x+1)(x-1)] = 2^3\]